Two years ago, two cats named Tony and Tiny, had 15 years of life together. Today, Tony is 13 years old. In how many years will Tiny be 9 years old?
Answer by lwsshak3(11628)

Two years ago, two cats named Tony and Tiny, had 15 years of life together. Today, Tony is 13 years old. In how many years will Tiny be 9 years old?
***
2 yrs ago:
let x=tony's age
15-x=tiny's age
..
at present:
x+2=tony's age=13
x=11
15-x+2=6=tiny's age
..
In how many years will Tiny be 9 years old? 3